OpenStudy (anonymous):

If a 14-foot-diameter sheave is used to drive a cable car, at what angular velocity must sheave turn in order for the cable car to travel 11 miles per hour? Give your answer in revolution per minute.(1 mi = 5280ft)

OpenStudy (anonymous):

radius=14/2=7 circumference C=2 pi r \[=2*\frac{ 22 }{ 7 }*7=44ft.\] no .of revolutions in 1 mile \[=\frac{ 5280 }{ 44 }=120\] no. of revolutions in 11 mile=120*11=1320 no. of revolutions in 1 hr.=1320 no. of revolutions per minute=1320/60=?

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

yw